OpenStudy (anonymous):

integrate 1/x^4-1 (using trig substitution)

OpenStudy (anonymous):

(x^4 - 1) = (x^2 - 1)(x^2 + 1) = (x - 1)(x + 1)(x^2 + 1) split it up into partial fractions...and you'll get arctan(x) and some logs...

OpenStudy (anonymous):

You have the form\[x ^{2}-a ^{2}\]\[a ^{2}\sec ^{2}\theta -a ^{2}\]\[a ^{2}\tan ^{2}\theta\]

OpenStudy (anonymous):

x=a sec theta

OpenStudy (anonymous):

In this case with x^4, x^2=a sec theta